# Towards Fast and Scalable Normal Integration using Continuous Components

### v1

@@ 4.3 Evaluation of scalability
Figure 4: Evaluation on large-scale normal maps. All experiments use $\Delta E_{\mathrm{max}}=10^{-3}$, $T=15$, outlier reweighting, and $\mathrm{freq}_{\mathrm{merging}}=5$ (where applicable). Total execution time $t$ and, where ground-truth reconstruction is available, mean relative depth error (RE) are reported. $\mathrm{M}$ and $\mathrm{NM}$ denote our version with and without merging, respectively. Sources of the normal maps: col. 11-22: rendered from meshes [9], [7]; col. 33-44: from photometric stereo [13], data from [13]; col. 55-66: predicted by DSINE [3] on in-the-wild images [3], [6].
@@
To assess the scalability of component-based and pixel-level formulations, we evaluate our method and the baseline on a set of mid-to-high-resolution normal maps that, unlike those from object-level datasets, have no masked-out pixels. We include normals rendered from meshes (using BlenderProc [8]), real-world normals from photometric stereo [13], and normals predicted by the state-of-the-art normal estimation method DSINE [3] on in-the-wild images. The number of pixels with a valid normal vector varies between 311 296311\,296 and 1 264 6401\,264\,640. We run our method with merging enabled (with $\mathrm{freq}_{\mathrm{merging}}=5$) both in its component-based version (with $\theta_{c}=2.0^{\circ}$) and in its pixel-level variant, including for the latter also the version without merging. For all methods, we use outlier reweighting, with convergence criteria $\Delta E_{\mathrm{max}}=10^{-3}$ and $T=15$.
@@
The results of our evaluation, together with the input normal maps and the output reconstructions, are shown in Fig.˜4. Our component-based framework achieves convergence on average one order of magnitude faster than our pixel-level version and the pixel-level method of [19]. This result is a consequence of the smaller scale of our optimization problem. Using our normal similarity criterion, our method is able to detect large continuous regions in the input maps and thereby reduce the number of optimization variables by a factor between $5$ and $50$. It follows that while pixel-level methods require several minutes for convergence even at the same resolution used in the DiLiGenT dataset (cf. first column in Fig.˜4), our method converges in a few seconds for the same input and in less than $2$ minutes for the larger resolutions. While the log-depth-based formulation of [19] converges on average faster than our pixel-level variant under the same conditions, we note that, as opposed to the small-scale DiLiGenT normal maps, enabling merging for our version on large-scale normal maps results in significantly reduced runtime with little to no impact on the reconstruction. This highlights that merging can be particularly beneficial in settings where iterations with the full number of pixels are costly and therefore a reduction in the number of variables is desirable.
@@ Appendix B Detailed profiling of our run time
Table 2: Profiling of our method on the the DiLiGenT benchmark [22]. Intermediate execution times from the start, after the completion of each step are reported $\mathrm{s}$]. $\theta_{c}=3.5^{\circ}$, $\Delta E_{\mathrm{max}}=10^{-3}$, $T=150$, 88-connectivity, and outlier reweighting are used, without merging.
@@ Appendix C Ablation on outlier reweighting
Table 4: Ablation on the outlier reweighting mechanism on the DiLiGenT benchmark [22]. The mean absolute depth error (MADE) $\mathrm{m}\mathrm{m}$] of our method is reported. The upper outlier reweighting threshold $U$ is set to $10^{-3}$, and soft threshold with different lower thresholds $L$ as well as hard thresholding based on $U$ are compared. All experiments use convergence criteria $\Delta E_{\mathrm{max}}=10^{-3}$ and $T=150$, 88- pixel connectivity, without merging.
@@ Appendix D Ablation on the convergence parameters
Table 5: Ablation on the convergence criteria $\Delta E_{\mathrm{max}}$ and $T$ on the DiLiGenT benchmark [22]. The mean absolute depth error (MADE, abbreviated as $\mathrm{Err}$) $\mathrm{m}\mathrm{m}$] and the total execution time (abbreviated as $t$) $\mathrm{s}$] of our method are reported. All experiments use outlier reweighting $W^{\mathrm{out}}_{b\rightarrow a}$ (20) with $L=10^{-5}$ and $U=10^{-3}$, without merging.
@@ Appendix E Ablation on the threshold for component formation
Figure 6: Continuous components identified by our method for different normal similarity thresholds $\theta_{c}$ and 88- connectivity, DiLiGenT benchmark [22]. For each object and threshold, different colors indicate different components. Below each component image are: the minimum mean average depth error (MADE, in $\mathrm{mm}$) that can be theoretically achieved by scaling the continuous components; the number of components.
@@ Appendix F Ablation on the effect of merging in the DiLiGenT benchmark
Figure 7: Continuous components at different stages $m$ of the merging process, DiLiGenT benchmark [22]. For each object and threshold, different colors indicate different components. Below each component image are: first row: the minimum mean average depth error (MADE, in $\mathrm{mm}$) that can be theoretically achieved by scaling the continuous components and the MADE ($\mathrm{m}\mathrm{m}$) achieved by the optimization at the end of the merging stage, the latter in brackets and underlined; second row: number of components. For each object, $M_{\mathrm{tot}}$ denotes the total number of merging operations required to obtain a single component. For all objects, normal similarity threshold $\theta_{c}=3.5^{\circ}$ and 88- connectivity are used, with $\mathrm{freq}_{\mathrm{merging}}=5$ and $\Delta E_{\mathrm{max}}=10^{-3}$.
@@ Appendix G Ablation on pixel connectivity
Table 6: Ablation on the pixel connectivity (abbreviated as $\mathrm{Conn.}$) on the DiLiGenT benchmark [22]. The mean absolute depth error (MADE, abbreviated as $\mathrm{Err}$) $\mathrm{m}\mathrm{m}$] and the total execution time (abbreviated as $t$) $\mathrm{s}$] of our method are reported. All experiments use outlier reweighting $W^{\mathrm{out}}_{b\rightarrow a}$ (20) with $L=10^{-5}$ and $U=10^{-3}$, convergence criteria $\Delta E_{\mathrm{max}}=10^{-3}$ and $T=150$, without merging.
@@ Appendix H Limitations
Figure 8: Example corner case not handled by our component-formation heuristic. When the input normals are continuous across a surface discontinuity, our heuristic for component formation cannot detect separate components. Since the model of discontinuity that we adopt [5] cannot recover discontinuities under the same corner case, the discontinuity will not be incorporated during the component filling and the two incorrectly merged pieces of surface will jointly adjust their scale in subsequent steps of the optimization. Source of the input normal map and ground-truth surface visualization: [5] (Fig. 14, Supplementary).
@@
Figure 9: Example limitation due to overly-smooth input normal map. In the case of overly-smooth input normal maps, our heuristic for component formation cannot detect separate components across the non-sharp boundaries, as is the case in this example of the boundary between the sticks in the foreground and the background. As a consequence, the foreground object is partly merged into the background. Source of input color image: [14].
Component formation and model of discontinuity. Since our heuristic for component formation relies on the similarity between neighboring normals, if the normals are continuous across a surface discontinuity it cannot detect disconnected surface regions as separate components, as depicted in the example of Fig.˜8. We note, however, that this example represents a corner case more in general of normal integration methods, as previously noted for instance by [5] (Fig. 14, Supplementary). In particular, in this case the integration problem itself is ill-posed, since in this setting it is not possible to determine whether a discontinuity is present from the normal map alone.
A similar issue arises in the case of overly-smooth input normals, with indistinct boundaries, which sometimes occur in normal maps produced as output by learning-based approaches for normal estimation, such as DSINE [3]. For these normals maps, our heuristic for component formation might sometimes merge multiple surfaces into a single component, depending on the similarity threshold $\theta_{c}$ (Fig.˜9). We note, however, that even for an incorrect component decomposition (i.e., that merges surface regions separated by a discontinuity) the corresponding reconstruction can in theory still correctly capture the discontinuity if the model of discontinuity is able to describe it. This is because in the initial phase each surface patch is reconstructed using the same model of discontinuity deployed for relative scale optimization. Viceversa, if the model of discontinuity is unable to capture the discontinuity, as is the case for full discontinuities such as the foreground-background boundaries in Fig.˜9, the general problem of retrieving the scale of surface regions across such discontinuities cannot be addressed without additional knowledge.

### v2

@@ 4.3 Evaluation of scalability
Figure 4: Evaluation on large-scale normal maps. All experiments use $\Delta E_{\mathrm{max}}=10^{-3}$, $T=15$, outlier reweighting, and $\mathrm{freq}_{\mathrm{merging}}=5$ (where applicable). Total execution time $t$ and, where ground-truth reconstruction is available, mean relative depth error (RE) are reported. $\mathrm{M}$ and $\mathrm{NM}$ denote our version with and without merging, respectively. Sources of the normal maps: col. 11-22: rendered from meshes [MeshCozyModernBedroom2023], [MeshCozyLivingRoomBaked2022]; col. 33-44: from photometric stereo [Ikehata2023SDMUniPS], data from [Ikehata2023SDMUniPS]; col. 55-66: predicted by DSINE [Bae2024DSINE] on in-the-wild images [Bae2024DSINE], [ImageWinterWeddingCake2007].
@@
To assess the scalability of component-based and pixel-level formulations, we evaluate our method and the baseline on a set of mid-to-high-resolution normal maps that, unlike those from object-level datasets, have no masked-out pixels. We include normals rendered from meshes (using BlenderProc [Denninger2023BlenderProc]), real-world normals from photometric stereo [Ikehata2023SDMUniPS], and normals predicted by the state-of-the-art normal estimation method DSINE [Bae2024DSINE] on in-the-wild images. The number of pixels with a valid normal vector varies between 311 296311\,296 and 1 264 6401\,264\,640. We run our method with merging enabled (with $\mathrm{freq}_{\mathrm{merging}}=5$) both in its component-based version (with $\theta_{c}=2.0^{\circ}$) and in its pixel-level variant, including for the latter also the version without merging. For all methods, we use outlier reweighting, with convergence criteria $\Delta E_{\mathrm{max}}=10^{-3}$ and $T=15$.
@@
The results of our evaluation, together with the input normal maps and the output reconstructions, are shown in Fig.˜4. Our component-based framework achieves convergence on average one order of magnitude faster than our pixel-level version and the pixel-level method of [Milano2025DiscontinuityAwareNormalIntegration]. This result is a consequence of the smaller scale of our optimization problem. Using our normal similarity criterion, our method is able to detect large continuous regions in the input maps and thereby reduce the number of optimization variables by a factor between $5$ and $50$. It follows that while pixel-level methods require several minutes for convergence even at the same resolution used in the DiLiGenT dataset (cf. first column in Fig.˜4), our method converges in a few seconds for the same input and in less than $2$ minutes for the larger resolutions. While the log-depth-based formulation of [Milano2025DiscontinuityAwareNormalIntegration] converges on average faster than our pixel-level variant under the same conditions, we note that, as opposed to the small-scale DiLiGenT normal maps, enabling merging for our version on large-scale normal maps results in significantly reduced runtime with little to no impact on the reconstruction. This highlights that merging can be particularly beneficial in settings where iterations with the full number of pixels are costly and therefore a reduction in the number of variables is desirable.
@@ Appendix B Detailed profiling of our run time
Table 2: Profiling of our method on the the DiLiGenT benchmark [Shi2016DiLiGenT]. Intermediate execution times from the start, after the completion of each step are reported $\mathrm{s}$]. $\theta_{c}=3.5^{\circ}$, $\Delta E_{\mathrm{max}}=10^{-3}$, $T=150$, 88-connectivity, and outlier reweighting are used, without merging.
@@ Appendix C Ablation on outlier reweighting
Table 4: Ablation on the outlier reweighting mechanism on the DiLiGenT benchmark [Shi2016DiLiGenT]. The mean absolute depth error (MADE) $\mathrm{m}\mathrm{m}$] of our method is reported. The upper outlier reweighting threshold $U$ is set to $10^{-3}$, and soft threshold with different lower thresholds $L$ as well as hard thresholding based on $U$ are compared. All experiments use convergence criteria $\Delta E_{\mathrm{max}}=10^{-3}$ and $T=150$, 88- pixel connectivity, without merging.
@@ Appendix D Ablation on the convergence parameters
Table 5: Ablation on the convergence criteria $\Delta E_{\mathrm{max}}$ and $T$ on the DiLiGenT benchmark [Shi2016DiLiGenT]. The mean absolute depth error (MADE, abbreviated as $\mathrm{Err}$) $\mathrm{m}\mathrm{m}$] and the total execution time (abbreviated as $t$) $\mathrm{s}$] of our method are reported. All experiments use outlier reweighting $W^{\mathrm{out}}_{b\rightarrow a}$ (20) with $L=10^{-5}$ and $U=10^{-3}$, without merging.
@@ Appendix E Ablation on the threshold for component formation
Figure 6: Continuous components identified by our method for different normal similarity thresholds $\theta_{c}$ and 88- connectivity, DiLiGenT benchmark [Shi2016DiLiGenT]. For each object and threshold, different colors indicate different components. Below each component image are: the minimum mean average depth error (MADE, in $\mathrm{mm}$) that can be theoretically achieved by scaling the continuous components; the number of components.
@@ Appendix F Ablation on the effect of merging in the DiLiGenT benchmark
Figure 7: Continuous components at different stages $m$ of the merging process, DiLiGenT benchmark [Shi2016DiLiGenT]. For each object and threshold, different colors indicate different components. Below each component image are: first row: the minimum mean average depth error (MADE, in $\mathrm{mm}$) that can be theoretically achieved by scaling the continuous components and the MADE ($\mathrm{m}\mathrm{m}$) achieved by the optimization at the end of the merging stage, the latter in brackets and underlined; second row: number of components. For each object, $M_{\mathrm{tot}}$ denotes the total number of merging operations required to obtain a single component. For all objects, normal similarity threshold $\theta_{c}=3.5^{\circ}$ and 88- connectivity are used, with $\mathrm{freq}_{\mathrm{merging}}=5$ and $\Delta E_{\mathrm{max}}=10^{-3}$.
@@ Appendix G Ablation on pixel connectivity
Table 6: Ablation on the pixel connectivity (abbreviated as $\mathrm{Conn.}$) on the DiLiGenT benchmark [Shi2016DiLiGenT]. The mean absolute depth error (MADE, abbreviated as $\mathrm{Err}$) $\mathrm{m}\mathrm{m}$] and the total execution time (abbreviated as $t$) $\mathrm{s}$] of our method are reported. All experiments use outlier reweighting $W^{\mathrm{out}}_{b\rightarrow a}$ (20) with $L=10^{-5}$ and $U=10^{-3}$, convergence criteria $\Delta E_{\mathrm{max}}=10^{-3}$ and $T=150$, without merging.
@@ Appendix H Limitations
Figure 8: Example corner case not handled by our component-formation heuristic. When the input normals are continuous across a surface discontinuity, our heuristic for component formation cannot detect separate components. Since the model of discontinuity that we adopt [Cao2022BiNI] cannot recover discontinuities under the same corner case, the discontinuity will not be incorporated during the component filling and the two incorrectly merged pieces of surface will jointly adjust their scale in subsequent steps of the optimization. Source of the input normal map and ground-truth surface visualization: [Cao2022BiNI] (Fig. 14, Supplementary).
@@
Figure 9: Example limitation due to overly-smooth input normal map. In the case of overly-smooth input normal maps, our heuristic for component formation cannot detect separate components across the non-sharp boundaries, as is the case in this example of the boundary between the sticks in the foreground and the background. As a consequence, the foreground object is partly merged into the background. Source of input color image: [ImageCakePops2013].
Component formation and model of discontinuity. Since our heuristic for component formation relies on the similarity between neighboring normals, if the normals are continuous across a surface discontinuity it cannot detect disconnected surface regions as separate components, as depicted in the example of Fig.˜8. We note, however, that this example represents a corner case more in general of normal integration methods, as previously noted for instance by [Cao2022BiNI] (Fig. 14, Supplementary). In particular, in this case the integration problem itself is ill-posed, since in this setting it is not possible to determine whether a discontinuity is present from the normal map alone.
A similar issue arises in the case of overly-smooth input normals, with indistinct boundaries, which sometimes occur in normal maps produced as output by learning-based approaches for normal estimation, such as DSINE [Bae2024DSINE]. For these normals maps, our heuristic for component formation might sometimes merge multiple surfaces into a single component, depending on the similarity threshold $\theta_{c}$ (Fig.˜9). We note, however, that even for an incorrect component decomposition (i.e., that merges surface regions separated by a discontinuity) the corresponding reconstruction can in theory still correctly capture the discontinuity if the model of discontinuity is able to describe it. This is because in the initial phase each surface patch is reconstructed using the same model of discontinuity deployed for relative scale optimization. Viceversa, if the model of discontinuity is unable to capture the discontinuity, as is the case for full discontinuities such as the foreground-background boundaries in Fig.˜9, the general problem of retrieving the scale of surface regions across such discontinuities cannot be addressed without additional knowledge.